50. Solve each system by substitution or addition, whichever is easier.
2y-x=3
x=3y-5
What are the steps to solving this problem?

Answer by doukungfoo(195)

2y-x=3
x=3y-5
Since the second equation has already been set equal to x I'll demostrate the substitution method.
Substitute 3y-5 for x in the first equation
2y - x = 3
2y - (3y-5) = 3
2y - 3y + 5 = 3
-1y + 5 = 3
-y = -2
y = 2
Now use this value for y to solve for x in the second equation
x = 3y - 5
x = 3(2) - 5
x = 6 - 5
x = 1
Check both anwsers in both equations
Try the first equation
2y - x = 3
2(2) - (1) = 3
4 - 1 = 3
3 = 3
Now try the second equation
x = 3y - 5
1 = 3(2) - 5
1 = 6 - 5
1 = 1